I need help with 2 questions 10y = 40 and 5 2/7 = ?/7

Answer by rothauserc(4718)

10y = 40
:
Divide both sides of = by 10
:
y = 4
:
5 2/7
:
There are (5 * 7) / 7 = (35 / 7) in 5, then
:
(35 / 7) + (2 / 7) = (37 / 7), therefore
:
5 2/7 = (35 / 7)
: